Fred’s Widget Company wants to make sure his widgets are within the weight tolerances he advertises on the radio. He says the weight of his widgets is 15 pounds. He samples 49 of his widgets and finds the mean weight of the widgets to be 14.5 with a standard deviation of 1.2 pounds. Based on this information, construct the 95 percent confidence interval for the sample.

a. What is the lower confidence limit? ____

b. What is the upper confidence limit? ____

Answer by stanbon(57222)

Fred’s Widget Company wants to make sure his widgets are within the weight tolerances he advertises on the radio. He says the weight of his widgets is 15 pounds. He samples 49 of his widgets and finds the mean weight of the widgets to be 14.5 with a standard deviation of 1.2 pounds. Based on this information, construct the 95 percent confidence interval for the sample.
-----
x-bar = 14.4
E = 1.96*(1.2/sqrt(49)) = 0.336
a. What is the lower confidence limit? 14.4-0.336 = 14.064____
b. What is the upper confidence limit? 14.4+0.336 = 14.736____
========================================
Cheers,
Stan H.